OpenStudy (anonymous):

can someone please help me? the problem is x/y+z/w=1+x. i keep getting (y(z-w))/(w(y+1))=x but when i checked it doesn't seem to be right. thanks :)

OpenStudy (anonymous):

what is the goal?

OpenStudy (anonymous):

oh sorry. I'm trying to find x

OpenStudy (anonymous):

$$\frac{x}y+\frac{z}w=1+x\\\frac{x}y-x=1-\frac{z}w\\\left(\frac1y-1\right)x=1-\frac{z}w\\x=\frac{1-z/w}{1/y-1}=\dots$$

OpenStudy (anonymous):

so$$x=\frac{1-z/w}{1/y-1}\cdot\frac{w}w=\frac{w-z}{w(1/y-1)}\cdot\frac{y}y=\frac{y(w-z)}{w(1-y)}$$

OpenStudy (anonymous):

why wouldn't you just multiply by wy to get rid of the fractions?

OpenStudy (anonymous):

that's what I just did